In a right triangle, a and b are the lengths of the legs and c is the length of the hypotenuse. If a = 70 millimeters and b = 24 millimeters, what is the perimeter? If necessary, round to the nearest tenth.

Answer by Cromlix(4381)

Pythagoras.
a^2 + b^2 = c^2
70^2 + 24^2 = c^2
c^2 = 5476
c = 74 mm
Therefore perimeter
= 70 + 24 + 74 = 168 mm
Hope this helps
:-)